OpenStudy (anonymous):

Describe how to estimate a non-perfect square root to the hundredths place without using a calculator.

OpenStudy (fibonaccichick666):

....evil, comes to mind. And anyways, you don't usually do more than say it is between two integers. There is no simple way to find it to the hundreths place that I know of.

OpenStudy (fibonaccichick666):

you do a bunch of guess and check

OpenStudy (amistre64):

theres a long hand divisiony process

OpenStudy (fibonaccichick666):

I would like to see that amistre, I don't actually know it.

OpenStudy (fibonaccichick666):

my way is just take a nice guess and compare, usually find it in 5 tries

OpenStudy (fibonaccichick666):

I mean unless you mean doing a prime factorization and multiplying by a common irrational root that you happen to know the digits of

OpenStudy (amistre64):

sqrt(4351.234) now the idea is that the biggest perfect square of a single digit is 9*9 = 81, 2 positions from the decimal start partitioning it into 2 digit parts sqrt(43 51. 23 40) now we are going to work off of (x+y)^2 = x^2 + 2xy + y^2 we want to find x^2 that divides 43 and subtract it .... like long division 6 ------------ sqrt(43 51. 23 40) 36 --- 7 51 <-- drop down now comes the 2xy part, we have x already given as 6 2*6 = 12 + y some single digit soo 6 y ------------ sqrt(43 51. 23 40) 36 --- 12+y| 7 51 (12+y)*y < 751

OpenStudy (anonymous):

Btw this is about estimating it not doing it on paper. i cant do that in my head.

OpenStudy (fibonaccichick666):

the problem says nothing about getting to use paper, just no calculator?

OpenStudy (fibonaccichick666):

also, very very interesting amistre! When did you learn that?

OpenStudy (anonymous):

The lesson is on estimating

OpenStudy (fibonaccichick666):

you can use paper to estimate

OpenStudy (fibonaccichick666):

ie. find square root of 167, well it's between 12 and 13, closer to thirteen, I'd bet like 12.9 ish but I can't do two places without testing it by writing it on paper

OpenStudy (fibonaccichick666):

estimating is not always the same as mental math

OpenStudy (amistre64):

y = 5 is just right 125 * 5 = 625 6 5 ------------ sqrt(43 51. 23 40) 36 --- 125| 7 51 625 ----- 126 23 2(65) = 130 130y < 12623 ; 9 works 1309*9 = 11781 6 5 . 9 ------------ sqrt(43 51. 23 40) 36 --- 7 51 625 ----- 126 23 11781 ------- 842 40 2(659) = 1318 1318y*y < 84240 ; 7?6? , 6 is good 13186*6 = 6 5 . 9 6 ------------ sqrt(43 51. 23 40) 36 --- 7 51 625 ----- 126 23 11781 ------- 842 40 79116 ------- 5124 dbl chk (65.96)^2 = 4350.7216

OpenStudy (amistre64):

when? i cant remember that far back lol

OpenStudy (fibonaccichick666):

lol, I would love to know why that works. I don't think I've ever seen it

OpenStudy (amistre64):

its based off of (x+y)^2 = x^2 + 2xy + y^2 finding xs and ys along the way the are "closest" and the rest is long divisiony

OpenStudy (fibonaccichick666):

But, how do you find s, I missed that part. (I don't really follow any of the logic though to be honest)

OpenStudy (amistre64):

when we divide long hand, whats the first thing we do? try to find a multiple that is closest to the digits we are assessing right?

OpenStudy (anonymous):



OpenStudy (fibonaccichick666):

yea, but I don't get why you split it up the way you did, I mean I know it has to be a two digit square, but I guess I just don't understand how that works.

OpenStudy (fibonaccichick666):

like could you prove that this method works for every number?

OpenStudy (amistre64):

9 is the largest digit, it makes the largest square, 9*9 = 81 so we need at least 2 digits of freedom to play with

OpenStudy (fibonaccichick666):

after partitioning, it would make sense that since your reasoning is the biggest 1 digit number squared is 2 digits, then why do we then assume a long division protocol for it. I don't get why it works, I want to know why

OpenStudy (amistre64):

ok start with x^2 we are finding x^2 that is closest to the required values right? then subtracting the difference

OpenStudy (amistre64):

(x+y)^2 = 1 42 x^2 + 2xy + y^2 = 1 42 x^2 + (2x + y)y = 1 42 (2x + y)y = 1 42 - x^2 but x^2 is a positional placement,

OpenStudy (fibonaccichick666):

But that doesn't explain why that works. You are performing a function not on the actual numbers as you are splitting it up, it becomes \(4300+51+.23+.004\) then finding the closest square to 4300. But I mean, how does that translate to (x+y)^2 then finding x and y

OpenStudy (amistre64):

x^2*100 is our positional value

OpenStudy (fibonaccichick666):

"sqrt(4351.234) now the idea is that the biggest perfect square of a single digit is 9*9 = 81, 2 positions from the decimal start partitioning it into 2 digit parts sqrt(43 51. 23 40) now we are going to work off of (x+y)^2 = x^2 + 2xy + y^2 " Why do we do this? Why does it work?

OpenStudy (amistre64):

10(2x) + 1y is our other positional setup, for the subsequent findings

OpenStudy (fibonaccichick666):

I don't understand what you are saying. I don't understand your terminology

OpenStudy (amistre64):

ugh ..... we dont have just "numbers" they are working on positions as well. as you noted: 42 isnt 42, its actually 4200 x^2 just gets us to 81 tops, x^2(100) is positionally correct

OpenStudy (fibonaccichick666):

I've never heard of positionally correct

OpenStudy (amistre64):

when we long hand divide: ------ 3 | 234 we find 7*3 = 21 but its not 7*3 we are finding, its 70*3